OpenStudy (anonymous):

The value of a dirt bike decreases by 25% each year. If you purchased this dirt bike today for $500, to the nearest dollar how much would the bike be worth 5 years later? $84 $119 $164 $222 is there a formula for this?

jimthompson5910 (jim_thompson5910):

The formula is F = P(1-r)^t where F is the future value, P is the present value, r is the rate of decrease (as a decimal) and t is the time in years

jimthompson5910 (jim_thompson5910):

In this case, P = 500, r = 0.25 and t = 5, so F = P(1-r)^t F = 500(1-0.25)^5

OpenStudy (anonymous):

let me do the math

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

118.65?

jimthompson5910 (jim_thompson5910):

good, now round to the nearest dollar

OpenStudy (anonymous):

$119

OpenStudy (anonymous):

thank man

jimthompson5910 (jim_thompson5910):

sure thing